Saturday 02/16/13

10:30 A – 8:30P, 10:00P – 11:55P

STORM Analysis

  • DaoSTORM analysis still running on Fab7-A647 Pc-750
  • Read write steps from USB connected hard-drive very slow when this analysis is running. (I’m not sure how much slower the analysis goes to have multiple files reading)
  • Really should upgrade this to eSATA or USB 3.0.
  • Setting up analysis of original Fab7 405-647 (01-31-13)
  • rerun z-calibration for orignal Fab7 data
  • Set up daoSTORM parameters and run on Tuck.
  • Setting up analysis of Fab7 multicolor data (with SS).
  • Movies 8-22 sample is outside of TIRF field. Data deleted. Should take more images! Signed up for STORM2 time tonight and tomorrow (Sunday) night.
  • modifying STORMfinder to use scratchPath (unecessary reading and writing from a disk which is being heavily accessed for other analysis is painfully slow).

Writing analysis script

  • BXC_s3_vs_s2_2012_02_16 – loop through all mlists, create images of all dots, compute statistics: 16x16nm pixels occupied by cluster, mean median max number of localizations per pixel bin, total localizations in cluster, total localization density (per spatial units of 16×16 bins).

Starting to explore:

BXC-S3cell_vs_S2Rcell (2 randomly selected examples, may not be at all representative of true data).

Sort which files have any analysis run

  • Don’t have DaoSTORM analysis results for BX-C. Running now on Monet
  • completed in 5 hrs.


  • Problem with color scaling is that GenGaussianSRImage converts to uint16 by rescaling from 0 to 2^16. This is fine for single Z but is a problem for Z as color.
  • solution — GenGaussianSRImage should return raw data as a single, this can be rescaled appropriately to max intensity of all pixels. Image of fixed solution: (previously all color scalebars went from 0 to 2^16: contrast_scalingZ

Fly work

  • MTD x PSc 2 from 2/1 and 1/28: definitely viable
  • BUT: MTD x Psc2 from 2/11 still 1 live male, 2 females 100+ embryos NOT viable !?
  • MAT alpha Tub x Psc2 (F1s): NOT viable!
  • expanding shRNA lines (need males). Set up new Psc2 x Mat alpha Tub crosses soon
  • New collection bottles: Ph-p,Ph-d/FM7 , Pc1/TM1, and SCM/TM3
  • Flipped all emerging F1s from MAT test crosses stored at 29C into new vials to check viability. Should get good confirmation results early next week.
  • Ph and Scm flies laying strong. Pc not laying well
  • PM virgin collection

Embryo work

  • Fixed Esc embryos, (7-20 hrs).
  • 11:55P flip collection cages
This entry was posted in Fly Work, Summaries and tagged , , . Bookmark the permalink.